Write the standard form of the equation of the line passing through (5, 2) and perpendicular to 2x-5y=5x+2y
Answer by lwsshak3(11628)

Write the standard form of the equation of the line passing through (5, 2) and perpendicular to 2x-5y=5x+2y
**
Equation of a straight line: y=mx+b, m=slope, b=y-intercept
2x-5y=5x+2y
7y=-3x
y=-3x/7
m=-3/7
..
For perpendicular line:
slope,m=7/3 (negative reciprocal of slope of given line)
equation:
y=7x/3+b
using given point (5,2), to solve for b
2=7*5/3+b
2=35/3+b
6=35+3b
3b=-29
b=-29/3
equation:
y=7x/3-29/3